2/3x + 3/4x = -17 Solve for the goven variable.
I'm woring on a test and it fuigures when I ask my teacher for help she shows me tutorials.. which none of the tutorials show fraction problems like this. Please show me how to do this. Thank you.

Answer by oberobic(2304)

2/3x + 3/4x = -17
.
One way to decide on the common denominator is to multiply them.
3*4 = 12.
So, you need to convert both fractions to 12ths.
.
2/3 * 4/4 = 8/12
3/4 * 3/3 = 9/12
.
8/12*x + 9/12*x = -17
.
17/12*x = -17
.
Multiply both sides by 12 to eliminate the fraction.
.
17x = -17*12
.
divide both sides by 17
.
x = -12
.
Done